# Nominal Evaluation of Automatic Multi-Sections Control Potential in Comparison To A Simpler One- or Two-Sections Alternative With Predictive Spray Switching

### v1

@@ Abstract
Automatic Section Control (ASC) is a long-standing trend for spraying in agriculture. It promises to minimise spray overlap areas. The core idea is to (i) switch off spray nozzles on areas that have already been sprayed, and (ii) to dynamically adjust nozzle flow rates along the boom bar that holds the spray nozzles when velocities of boom sections vary during turn maneuvers. ASC is not possible without sensors, in particular for accurate positioning data. Spraying and the movement of modern wide boom bars are highly dynamic processes. In addition, many uncertainty factors have an effect such as cross wind drift, boom height, nozzle clogging in open-field conditions, and so forth. In view of this complexity, the natural question arises if a simpler alternative exist. Therefore, an Automatic Multi-Sections Control method is compared to a proposed simpler one- or two-sections alternative that uses predictive spray switching. The comparison is provided under nominal conditions. Agricultural spraying is intrinsically linked to area coverage path planning and spray switching logic. Combinations of two area coverage path planning and switching logics as well as three sections-setups are compared. The three sections-setups differ by controlling 48 sections, 2 sections or controlling all nozzles uniformly with the same control signal as one single section. Methods are evaluated on 10 diverse real-world field examples, including non-convex field contours, freeform mainfield lanes and multiple obstacle areas. A preferred method is suggested that (i) minimises area coverage pathlength, (ii) offers intermediate overlap, (iii) is suitable for manual driving by following a pre-planned predictive spray switching logic for an area coverage path plan, and (iv) and in contrast to ASC can be implemented sensor-free and therefore at low cost.
@@ 1 Introduction
Agricultural spraying is also intrinsically linked to the development of logics for spray switching, i.e., for the on and off switching of nozzles. Besides optimised area coverage path planning, a second strategy for efficient spraying and saving of resources is to suitably switch on and off individual nozzles for areas already sprayed, and to vary nozzle flow rates of individual nozzles along turn maneuvers to compensate for different traveling velocities of different sections of the boom bar. The name for the second strategy is Automatic Section Control (ASC) or Automatic Boom Section Control [6]. ASC is a long-standing trend for spray applications in agriculture [7].
@@
Fig. 1: Illustration of switching results stemming from a map-based Automatic Section Control. The figure is from [21]. Four comments are made: (i) The gaps between green areas are due to the fact that boom section states (nozzle on or off states) are sampled discretely. (ii) Multiple overlapping green areas can still be observed. (iii) The discrete nature of the displayed results (large gaps between two sampled green areas) implies that delays between switching changes (from nozzle on- to an off-state and vice versa) naturally occur. (iv) It can further be observed that adjacent path segments (mainfield lanes) do not have a constant distance to each other. This results in excessive overlapping potential for the given example. The figure is given for visualisation of 4 typical difficulties associated with ASC.
This raises the natural question: ”In view of the plethora of complexities of the spraying process, is ASC worth it and does an alternative exist?
@@
For comparison, an alternative practical method is suggested that varies from a common ASC-solution along 2 independent vectors. First, the predictive spray switching method from [22] is employed. It is here extended to the non-convex general case with non-convex field areas and multiple obstacle areas such as tree islands, ponds, or power pole masts inside the field area. This method employs a predictive switching logic tailored to a specific path pattern. This is in contrast to the reactive switching logic that is common for ASC, which is based on Boustrophedon-based path planning (Ancient Greek for ”like the ox turns”) with a meandering path pattern.
Second, three different setups for the number of sections are considered: 1, 2 or multiple sections. See Fig. 2 for illustration. It is noted that in all three cases always the same number of nozzles is controlled. However, nozzles are then controlled either individually or uniformly. For example, for the 1-section setup all nozzles along the boom bar uniformly receive the same switch-on or switch-off signal. In contrast, for the multiple sections-setup different sections can receive different switch-on or switch-off signals.
@@
Thus, in total 6 different experimental setups are compared. Evaluation metrics are (i) pathlength, (ii) spray volume consumption in litres (a proxy for the overlap area, but more accurate since also accounting for nozzle flow rate variations for turn compensation), and (iii) overlap area visualisation for qualitative verification.
The remaining paper is organised as follows: materials and methods, results, discussion and the conclusion are described in Sections 2-5.
@@ 19 Pulse-Width-Modulation (PWM) for [17, 7]
Fig. 3: Complexities of the spraying process: a sketch illustrating different velocities along different sections of the boom bar during turn maneuvers. Rotational velocity is denoted by $\dot{\psi}$. The distance of a boom section $i$ from the center of rotation is indicated by $l_{i}$.
Fig. 4: Complexities of the spraying process: a sketch illustrating (i) boom height $H$, (ii) inter-nozzle spacing $w_{\text{nozzle}}$, (iii) spray angle $\gamma$, (iv) angle $\eta$ for angled nozzles along the vertical axis (’z-axis’) such that adjacent spray patterns do not intersect, and (v) typically ellipsoid spray patterns for flat-fan nozzles.
@@
Table 1 lists factors that influence uncertainty of the spraying process, particularly in open-space agriculture. A list is provided for clarity and to underline the complexity of the spraying process, with 20 factors that all affect the spraying result.
@@ 2.2 Four Hierarchical Planning Levels
Fig. 6: Distinction between 4 hierarchical planning levels for area coverage in agriculture: (i) path planning-level, (ii) block switching-level, (iii) sections switching-level, and (iv) a nozzle control-level (realisation of individual nozzle actuation commands at the embedded machine level). The topic of this paper are the first three levels (i)-(iii).
@@
Note that the spraying process can be further facilitated by the design of the seeding process, which precedes the spraying process. Two influences are mentioned. First, the orientation of straight mainfield lanes or alternatively freeform mainfield lanes, for example, adapted in shape using a partial segment of the headland path as reference, can facilitate turning or reduce the required number of transitions between mainfield lanes and headland path [27]. Second, the intersection line between mainfield area and headland area can be used as visual cue to determine suitable spray on/off-switching locations. See Fig. 7.
@@
Fig. 7: Illustration of steps from field contour data to area coverage. The practical usefulness of different planting directions in headland and mainfield area is highlighted. The intersection line between the two areas is useful for triggering spray switching changes (i.e., on- and off-switching) and is useful for both, (i) the Boustrophedon-based and (ii) the Alternative planning method. An extension to the case with obstacle areas (tree islands) inside the field is straightforward by the introduction of obstacle area headland areas with planting areas different from the mainfield area. The headland and mainfield planting directions in above figure are sketched at a $90^{\circ}$-angle to each other. Obviously, any other angle is also possible as long as an intersection line results that can be employed as a visual cue for switching commands.
@@ 2.3 Area Coverage Path Planning Level
Two different area coverage path planning strategies are compared. For convexly-shaped fields these can be used throughout. For the non-convex cases a graph optimisation problem is solved that adopts the two respective strategies whenever possible.
@@ 2.4 Block Switching Level
According to the hierarchical structure of Fig. 6, a block swiching logic is applied on top of a given area coverage path plan according to the methods from the previous Sect. 2.3.
@@ 2.5 Sections Switching Level
According to the hierarchical structure of Fig. 6, on the sections switching level (i) a decision regarding section width, and (ii) a decision regarding discretisation modeling has to be made. There are 2 factors that influence our discretisation modeling of spray section areas at each sampling distance along a path.
The 2 influencing factors are: (i) Spray patterns that result from the dominant standard flat-fan nozzle type are ellipsoid [13]. (ii) The simplifying assumption is made that the boom bar is orthogonally oriented along the area coverage path plan at every sampling point. See Fig 4 for the visualisation of ellipsoidal spray patterns. See Fig. 14 for illustration of assumption (ii).
@@
The second polygon type is constructed at heading changes along the area coverage path and constructed by connecting boundary points of two polygons at the previous and next sampling space. At those heading changes the traveling velocity along different sections of the boom bar varies according to the formula
@@
where $N>0$ different sections are assumed both at the left and right hand-side of the machinery path; see Fig. 14 for illustration. Depending on a left-turn or a right-turn the traveling velocity of a boom section may be larger or smaller than the machinery reference velocity $v_{\text{ref}}$. Negative $v_{i}<0$ are possible, which imply ’backwards motion’ of section $i$ that can occur during turn maneuvers and low reference velocity $v_{\text{ref}}\approx 0$.
The assumption is made that spray volumes are uniform over the polygon areas for this discretisation scheme.
A section area width $w>0$ has to be selected. This is not a straightforward choice since lateral transients result from spray overlap as Fig. 5 illustrated. The smallest possible selection is
@@
where $w_{\text{nozzle}}>0$ indicates the inter-nozzle spacing as illustrated in Fig. 4. For the evaluation of numerical experiments below this setting is used since it offers the finest discretisation spacing possible.
Nevertheless, the simplification of assuming constant spray along this section despite an actual lateral transient according to Fig. 5 is explicitly stressed. In Fig. 5, three nozzles are illustratively treated as one section of width $w_{\text{section}}>0$ with 100% overlap, i.e., the lateral spray distance on the ground is twice the nozzle spacing [8].
We assume the following linear relationship [8, 13] between controllable section flow rates, $f_{\text{section}}^{i}$ [l/s] for $N$ sections $i\in\{1,\dots,N\}$ on both the left- and right-hand side of the machinery path according to Fig. 4, and actual spray volume $s_{\text{volume}}^{\text{actual},i}$ $\ell$/ha],
@@
where $w>0$ denotes the section width and $v_{i}\in\mathbb{R}$ denotes the traveling speed of the centroid of section $i$ according to (1). Throughout visualisations in this paper and in line with Fig. 14 and (3), smaller and larger $s_{\text{volume}}^{\text{actual},i}$ are differentiated by brighter and darker gray colors, respectively.
Ideally, for uniform spraying of the entire field area the applied spray volume is constant for all non-overlapping section areas with
@@
Second, one may switch off, $f_{\text{section}}^{i}=0$, if there is an overlap with an already sprayed area or with an area that will be sprayed later within an predictive spraying scheme. The standard approach for this is an occupancy grid-based approach [23, 21]. The advantage of this approach is computational simplicity. The disadvantage is that at least 1 hyperparameter needs to be set, which results in either spray gaps or spray overlaps, as Fig. 15 demonstrates. Therefore, an alternative polygon-based approach is employed throughout the remainder of this paper: Here, (i) a polygon is recursively constructed from the area that has been sprayed so far, before (ii) at area coverage runtime for a potential spray section cell candidate it is evaluated whether it is element of the polygon representing already sprayed area or whether it is not. If the centroid of a cell candidate is not yet element of the polygon the cell is sprayed and the polygon is augmented. For computational efficiency and to keep a constant size, the polygon is recursively updated by considering only the last 10 sampling spaces. This polygon-approach is only applied along the headland path area.
@@
Fourth and as a detail, for nominal numerical experiments in Sect. 3, equation (5) is recalculated at each sampling space along the area coverage path. No rate constraints on $f_{\text{section}}^{i}$ are considered. This assumption is in line with the nominal experimental setup and best-case evaluation. Rate-constraints are regarded as PWM-associated uncertainties in Table 1.
@@
Fig. 15: The machinery path to be followed from the top-left to the rights is in blue. Sprayed area is in gray. Illustration of the disadvantage of an occupancy grid-based approach to determine if a small area has already been sprayed before or not (cf. also Fig. 1). The effect of different hyperparameter choices $d_{\text{G}}$, a threshold that determines if the centroid of a small area candidate is close to the centroid of an already sprayed area via the Euclidean norm, is visualised: In (a) there are gaps if $d_{\text{G}}$ is too large. In (b)-(c) there are overlaps if $d_{\text{G}}$ is too small. In contrast, in Fig. 15(d) an alternative non-occupancy grid based approach is visualised. This is approach does not require a threshold hyperparameter and yields both less gaps and less overlapping.
@@
Fig. 16: The effect of mainfield lane endings not being rectangular with respect to the headland path is visualised for the 3 different section-setups according to Fig. 2. See particularly the differences around the transition annotated with index number 6: In Fig. 16(a) all nozzles are simultaneously switched on or off. In contrast, in Fig. 16(c) each nozzle is switched individually. For the case of a single section, where all nozzles uniformly follow the same on/off-switching control signal, significant overlapping can result.
@@ 2.6 Strategy 1: Automatic Section Control
First, in contrast to an occupancy grid approach that was mentioned in Sect. 2.5 and that grids a field area into cells (i.e. discretised sub-areas of the field), here a different approach is taken. This is done to generate high-precision results and to avoid having to select discretisation-hyperparameters (see Fig. 15). Here the approach is used that (i) recursively extends a polygon that is constructed from the sprayed area so far, before (ii) evaluating at area coverage runtime if a potential spray section area is element of the polygon area or not.
@@
Finally, two additional observations are made. First, ASC must run automatedly since a large number of individually controlled nozzles (e.g. 48 nozzles for a 24m wide boom bar and 0.5m nozzle spacing) cannot be efficiently switched manually by a human operator at typical fast machinery traveling speeds. Second, at least one localisation sensor (e.g. for accurate RTK-GPS) is required to enable ASC-automation. Ideally, additional sensors are available to as accurately as possible measure the traveling velocity of individual nozzles (instead of estimating and inferring them from geometrical calculations and uncertainties such as mechanical oscillations that may occur for a wide boom bar during turn maneuvers). Thus, in general accurate ASC is a method that is highly dependent on accurate sensor measurements and automation.

### v2

@@ Abstract
Automatic Section Control (ASC) is a long-standing trend for spraying in agriculture. It promises to minimise spray overlap areas. The core idea is to (i) switch off spray nozzles on areas that have already been sprayed, and (ii) to dynamically adjust nozzle flow rates along the boom bar that holds the spray nozzles when velocities of boom sections vary during turn maneuvers. ASC is not possible without sensors, in particular for accurate positioning data. Spraying and the movement of modern wide boom bars are highly dynamic processes. In addition, many uncertainty factors have an effect such as cross wind drift, boom height, nozzle clogging in open-field conditions, and so forth. In view of this complexity, the natural question arises if a simpler alternative exist. Therefore, an Automatic Multi-Sections Control method is compared to a proposed simpler one- or two-sections alternative that uses predictive spray switching. The comparison is provided under nominal conditions. Agricultural spraying is intrinsically linked to area coverage path planning and spray switching logic. Combinations of two area coverage path planning and switching logics as well as three sections-setups are compared. The three sections-setups differ by controlling 48 sections, 2 sections or controlling all nozzles uniformly with the same control signal as one single section. Methods are evaluated on 10 diverse real-world field examples, including non-convex field contours, freeform mainfield lanes and multiple obstacle areas. An economic cost analysis is provided to compare the methods. A preferred method is suggested that (i) minimises area coverage pathlength, (ii) offers intermediate overlap, (iii) is suitable for manual driving by following a pre-planned predictive spray switching logic for an area coverage path plan, and (iv) and in contrast to ASC can be implemented sensor-free and therefore at low cost. Surprisingly strong economic arguments are found to not recommend ASC for small farms.
@@ 1 Introduction
$A_{\text{total}}$ Total farm size, (ha).
[…]
$\Delta K_{\text{ASC}}$ Purchase price difference for ASC, (EUR).
@@
$N_{\text{runs}}^{\text{field}}$ Number of field runs per crop cycle, (-).
$N_{\text{years}}^{\star}$ Number of years until profitability for ASC, (-).
@@
Agricultural spraying is also intrinsically linked to the development of logics for spray switching, i.e., for the on and off switching of nozzles. Besides optimised area coverage path planning, a second strategy for efficient and resource-saving spraying is to suitably switch on and off individual nozzles for areas already sprayed, and to vary nozzle flow rates of individual nozzles along turn maneuvers to compensate for different traveling velocities of different sections of the boom bar. The name for this is Automatic Section Control (ASC) or Automatic Boom Section Control [6]. ASC is a long-standing trend for spray applications in agriculture [7].
@@
Fig. 1: Illustration of switching results stemming from a map-based Automatic Section Control. The figure is from [21]. Four comments are made: (i) The gaps between green areas are due to the fact that boom section states (nozzle on- or off-states) are sampled discretely. (ii) Multiple overlapping green areas can still be observed. (iii) The discrete nature of the displayed results (large gaps between two sampled green areas) implies that delays between switching changes (from nozzle on- to an off-state and vice versa) naturally occur. (iv) It can further be observed that adjacent path segments (mainfield lanes) do not have a constant distance to each other. This results in excessive overlapping potential. The figure is given for visualisation of 4 typical difficulties associated with ASC.
This raises the natural question: ”In view of the plethora of complexities of the spraying process, is ASC quantitatively worth it and does an alternative exist?
@@
For comparison, an alternative practical method is suggested that varies from a common ASC-solution along 2 independent vectors. First, the predictive spray switching method from [22] is employed. It is here extended to the non-convex general case with non-convex field areas and multiple obstacle areas. This method employs a predictive switching logic tailored to a specific path pattern. This is in contrast to the reactive switching logic that is common for ASC, which is based on Boustrophedon-based path planning (Ancient Greek for ”like the ox turns”) with a meandering path pattern.
Second, three different setups for the number of sections are considered: 1, 2 or multiple sections. See Fig. 2 for illustration. In all three cases always the same number of nozzles is controlled. However, nozzles are controlled either individually or uniformly. For the 1-section setup all nozzles along the boom bar uniformly receive the same switch-on or switch-off signal. In contrast, for the multiple sections-setup different sections can receive different switch-on or switch-off signals.
@@
Thus, in total 6 different experimental setups are compared. Evaluation metrics are (i) pathlength, (ii) spray volume consumption in litres (a proxy for the overlap area, but more accurate since also accounting for nozzle flow rate variations for turn compensation), and (iii) overlap area visualisation. In addition, a detailed economic cost discussion is provided.
Materials and methods, results, discussion and the conclusion are described in Sec. 2-5.
@@ 19 Pulse-Width-Modulation (PWM) for [17, 7]
Fig. 3: Complexities of the spraying process: a sketch illustrating different velocities along different sections of the boom bar during turn maneuvers.
Fig. 4: Complexities of the spraying process: sketch illustrating (i) boom height $H$, (ii) inter-nozzle spacing $w_{\text{nozzle}}$, (iii) spray angle $\gamma$, (iv) angle $\eta$ for angled nozzles along the vertical axis (’z-axis’) such that adjacent spray patterns do not intersect, and (v) typically ellipsoid spray patterns for flat-fan nozzles.
@@
Table 1 lists factors that influence uncertainty of the spraying process, particularly in open-space agriculture. A list is provided for clarity and to underline complexity.
@@ 2.2 Four Hierarchical Planning Levels
Fig. 6: Distinction between 4 hierarchical planning levels for area coverage: (i) path planning-level, (ii) block switching-level, (iii) sections switching-level, and (iv) a nozzle control-level (realisation of individual nozzle actuation commands at the embedded machine level). The topic of this paper are the first three levels (i)-(iii).
@@
Note that the spraying process can be further facilitated by the design of the seeding process, which precedes the spraying process. Two influences are mentioned. First, the orientation of straight mainfield lanes or alternatively freeform mainfield lanes can facilitate turning or reduce the required number of transitions between mainfield lanes and headland path [27]. Second, the intersection line between mainfield area and headland area can be used as visual cue to determine suitable spray on/off-switching locations. See Fig. 7.
@@
Fig. 7: Illustration of steps from field contour data to area coverage. The practical usefulness of different planting directions in headland and mainfield area is highlighted. The intersection line between the two areas is useful for triggering spray switching changes (i.e., on- and off-switching) and is useful for both, (i) the Boustrophedon-based and (ii) the Alternative planning method. An extension to the case with obstacle areas (tree islands) inside the field is straightforward by the introduction of obstacle area headland areas with planting areas different from the mainfield area. The headland and mainfield planting directions in above figure are sketched at a $90^{\circ}$-angle to each other. Any other angle is also possible as long as an intersection line results that can be employed as a visual cue for switching command changes.
@@ 2.3 Area Coverage Path Planning Level
Two different area coverage path planning strategies are compared. For convexly-shaped fields these can be used throughout. For the non-convex cases a graph optimisation problem is solved that adopts the respective strategies whenever possible.
@@ 2.4 Block Switching Level
According to the hierarchical structure of Fig. 6, a block swiching logic is applied on top of a given area coverage path plan according to the methods from the previous Sec. 2.3.
@@ 2.5 Sections Switching Level
According to the hierarchical structure of Fig. 6, on the sections switching level (i) a decision regarding section width, and (ii) regarding discretisation modeling has to be made. There are 2 factors that influence our discretisation modeling of spray section areas at each sampling distance along a path.
The 2 influencing factors are: (i) Spray patterns that result from the dominant standard flat-fan nozzle type are ellipsoid [13]. (ii) The simplifying assumption is made that the boom bar is orthogonally oriented along the area coverage path plan at every sampling point. See Fig 4 for visualisation of ellipsoidal spray patterns. See Fig. 14 for illustration of assumption (ii).
@@
The second polygon type is constructed at heading changes along the area coverage path and constructed by connecting boundary points of two polygons at the previous and next sampling space. At those heading changes the traveling velocity along different sections of the boom bar varies according to
@@
where $N$ different sections are assumed both at the left and right hand-side of the machinery path; see Fig. 14 for illustration. Depending on a left-turn or a right-turn the traveling velocity of a boom section may be larger or smaller than the machinery reference velocity $v_{\text{ref}}$. Negative $v_{i}<0$ are possible, which imply ’backwards motion’ of section $i$ that can occur during turn maneuvers and at low reference velocity $v_{\text{ref}}\approx 0$.
Spray volumes are assumed to be uniform over the polygon areas for this discretisation scheme.
A section area width $w$ has to be selected. This is not a straightforward choice since lateral transients result from spray overlap (see Fig. 5). The smallest possible selection is
@@
where $w_{\text{nozzle}}$ indicates the inter-nozzle spacing as illustrated in Fig. 4. For numerical experiments below this setting is used since it offers the finest discretisation spacing possible.
Nevertheless, the simplification of assuming constant spray along this section despite an actual lateral transient according to Fig. 5 is explicitly stressed. In Fig. 5, three nozzles are illustratively treated as one section of width $w_{\text{section}}$ with 100% overlap, where the lateral spray distance on the ground is twice the nozzle spacing [8].
We assume the following linear relationship [8, 13] between controllable section flow rates, $f_{\text{section}}^{i}$ ($\ell$/s) for $N$ sections $i\in\{1,\dots,N\}$ on both the left- and right-hand side of the machinery path according to Fig. 4, and actual spray volume $s_{\text{volume}}^{\text{actual},i}$ ($\ell$/ha),
@@
where $w$ denotes the section width and $v_{i}\in\mathbb{R}$ denotes the traveling speed of the centroid of section $i$ according to (1). Throughout visualisations in this paper and in line with Fig. 14 and (3), smaller and larger $s_{\text{volume}}^{\text{actual},i}$ are differentiated by brighter and darker gray colors, respectively.
Ideally, for uniform spraying the applied spray volume is constant for all non-overlapping section areas with
@@
Second, one may switch off, $f_{\text{section}}^{i}=0$, if there is an overlap with an already sprayed area or with an area that will be sprayed later within an predictive spraying scheme. The standard approach for this is an occupancy grid-based approach [23, 21]. The advantage of this approach is computational simplicity. The disadvantage is that at least 1 hyperparameter needs to be set, which results in either spray gaps or spray overlaps, as Fig. 15 demonstrates. Therefore, an alternative polygon-based approach is employed throughout the remainder of this paper: Here, (i) a polygon is recursively constructed from the area that has been sprayed so far, before (ii) at area coverage runtime for a potential spray section cell candidate it is evaluated whether it is element of the polygon representing already sprayed area or not. If the centroid of a cell candidate is not yet element of the polygon the cell is sprayed and the polygon is augmented. For computational efficiency and to keep a constant size, the polygon is recursively updated by considering only the last 10 sampling spaces. This polygon-approach is only applied along the headland path area.
@@
Fourth and as a detail, for nominal numerical experiments in Sec. 3, equation (5) is recalculated at each sampling space along the area coverage path. No rate constraints on $f_{\text{section}}^{i}$ are considered. This assumption is in line with the nominal experimental setup and best-case evaluation. Rate-constraints are regarded as PWM-associated uncertainties in Table 1.
@@
Fig. 15: The machinery path to be followed from the top-left to the right is in blue. Sprayed area is in gray. Illustration of the disadvantage of an occupancy grid-based approach to determine if a small area has already been sprayed before or not (see also Fig. 1). The effect of different hyperparameter choices for $d_{\text{G}}$, a threshold that determines if the centroid of a small area candidate is close to the centroid of an already sprayed area via the Euclidean norm, is visualised. In (a) there are gaps if $d_{\text{G}}$ is too large. In (b)-(c) there are overlaps if $d_{\text{G}}$ is too small. In contrast, in Fig. 15(d) an alternative non-occupancy grid based approach is visualised. This is approach does not require a threshold hyperparameter and yields both less gaps as well as less overlapping.
@@
Fig. 16: The effect of mainfield lane endings not being rectangular with respect to the headland path is visualised for the 3 different section-setups according to Fig. 2. See in particular the differences around the transition annotated with index number 6: In Fig. 16(a) all nozzles are simultaneously switched on or off. In contrast, in Fig. 16(c) each nozzle is switched individually.
@@ 2.6 Strategy 1: Automatic Section Control
First, in contrast to an occupancy grid approach that was mentioned in Sec. 2.5 and that grids a field area into cells (i.e. discretised sub-areas of the field), here a different approach is taken. This is done to generate high-precision results and to avoid having to select discretisation-hyperparameters (see Fig. 15). Here the approach is used that (i) recursively extends a polygon that is constructed from the sprayed area so far, before (ii) evaluating at area coverage runtime if a potential spray section area is element of the polygon area or not.
@@
Finally, two additional observations are made: (i) ASC must run automatedly since a large number of individually controlled nozzles (e.g. 48 nozzles for a 24m wide boom bar and 0.5m nozzle spacing) cannot be efficiently switched manually by a human operator at typical fast machinery traveling speeds. (ii) At least one localisation sensor (e.g. for accurate RTK-GPS) is required to enable ASC-automation. Ideally, additional sensors are available to as accurately as possible measure the traveling velocity of individual nozzles (instead of estimating them from geometrical calculations subject to uncertainties such as mechanical oscillations that may occur for a wide boom bar during turn maneuvers). Thus, in general ASC is highly dependent on accurate sensor measurements and automation.